OpenStudy (anonymous):

find the dimension of the rectangle wiht a perimeter of 100 m that has the largest possible area.

OpenStudy (anonymous):

The maximum is \[A = l \times w\] the constraint is \[2l + 2w = 100\] in terms of x \[2x +2w = 100\] \[2w = 100 - 2x\] \[w = 50 - x\] A = x(50-x)

OpenStudy (anonymous):

A= 50x -x^2 end points (0, 50] find the first derivative A' = 50 - 2x set equal to zero x = 25

OpenStudy (anonymous):

x = 0, A = 0 x =25, A = 625 x =50 A = 0 so what's my dimension?